For a generic A→B, what quantity, when graphed versus time, will yield a straight line?
Match the words in the left column to the appropriate blanks in the sentences on the right.

A graph of -------- versus time yields a straight line for a second-order reaction. versus time yields a straight line for a second-order reaction.
1
Expert's answer
2020-03-23T10:15:17-0400

The second order reaction is given by "\frac{-dA}{dt}=k[A]^2"

"\frac{-dA}{[A]^2}=kdt"

integrating we get

"\frac{1}{[A]}-\frac{1}{[A]_o}=kt"

Hence graph of inverse of concentration at time t vs time t will give the straight line